# Adversarial Attacks of Vision Tasks in the Past 10 Years: A Survey

### v1

@@ 7.3.3. Metrics
Here, represents the set of attack samples, and is a binary function used to determine whether an attack sample is effective. Since the output of an LVLM is natural text, the criteria for the function can vary across different attack scenarios. For example, in image captioning tasks, the determination of effectiveness is often made by checking whether the output text diverges from the original meanings (untargeted) (Dong et al., 2023a) or close to the target semantics (targeted) (Wang et al., 2023b). In prompt injection and jailbreak attacks, success is usually determined by scrutinizing whether the output contains malicious contents (Lu et al., 2024; Fu et al., 2023) or violates usage policies (Chen et al., 2024a; Ying et al., 2024a). As summarized in Fig. 7, the existing implementations of the function in research can be categorized into four types:
@@
String Matching. This method checks whether the LVLM output contains certain keywords/phrases in the focus list or exactly matches a predefined target text (Lu et al., 2024; Fu et al., 2023). It is the most convenient strategy but can lead to false positives or negatives in some cases (e.g., when certain keywords are not in the focus list or the list is incomplete). In addition, the method is also known as Contain and ExactMatch (Liu et al., 2024d).
@@ 7.4.1. Attacker Knowledge
As shown in Fig. 8 (a), in LVLM adversarial attacks, we follow traditional criteria and categorize knowledge access into white-box, gray-box, and black-box. Unlike previous categories, we place query-based methods that require interaction with the victim model under the gray-box category, as these methods gather some degree of information in queries (e.g., estimating gradients in the text feature domain by LVLM’s responses (Zhao et al., 2024) or to generate/update attack prompts (Qraitem et al., 2024; Wu et al., 2023; Ma et al., 2024)), making them not entirely ignorant of the victims. Black-box methods refer to transfer-based attacks, which do not seek any information from the victims. These methods rely solely on prior knowledge, such as publicly available information and educated guesses. section 4 discusses the detailed classification criteria.
@@
Specifically, QT-Atk (Zhao et al., 2024) draws on the cascading approach of traditional transfer and query-based attacks (Cheng et al., 2019; Dong et al., 2021). It first applies transfer-based attacks to compute preliminary image perturbations and then refines the results (with PGD (Madry, 2017)) by estimating gradients using finite differences based on the model’s responses. All of (Qraitem et al., 2024; Wu et al., 2023; Ma et al., 2024) adopt the self-generate approach: the difference is that SGTA (Qraitem et al., 2024) requires identifying the most confusing class and description for a given image, while SASP (Wu et al., 2023) directly updates the attack prompt and VRP (Ma et al., 2024) incorporates role-playing dialogue. InstructTA (Wang et al., 2023b), on the other hand, assumes access to the victim’s visual encoder and applies a feature attack on it.
@@
Specifically, HADES (Li et al., 2024c) employs a three-stage strategy, progressively adding harmful information through typography, CIG, and adversarial perturbations to the attack samples. Focusing on disrupting LVLM’s CoT (Wei et al., 2022) process, SRA (Wang et al., 2024b) attacks both the reasoning and answer generation components. BMA (Bailey et al., 2023), JBA (Qi et al., 2024), UMK (Wang et al., 2024e), and BAP (Ying et al., 2024b) aggregate perturbations across datasets to create cross-prompt (Bailey et al., 2023)/corpus (Qi et al., 2024; Wang et al., 2024e; Ying et al., 2024b) universal adversarial images. AnyDoor (Lu et al., 2024) takes this further by binding universal adversarial images to specific text triggers (e.g., “SUDO”) to execute backdoor attacks. Unlike previously aggregated universal samples, CroPA (Luo et al., 2024a) achieves cross-prompt effects by continuously perturbing input text during sample “training”, which achieves aggregation over perturbed input text.
@@
Leveraging techniques such as typography and CIG, combined with specially designed text, to create attacks specific to LVLMs (Greshake et al., 2023; Gong et al., 2023b; Chen et al., 2023b; Shayegani et al., 2023; Liu et al., 2023d). (This type of methods is similar to that of white-box attacks, but reserchers report the cross-model transferability on it.)
@@
Specifically, CWA (Chen et al., 2023c) and SSA-CWA (Dong et al., 2023a) directly apply traditional transfer attack methods to test LVLMs, while ImgJP (Niu et al., 2024), AAMA (Wu et al., 2024), and OOD-VQA (Tu et al., 2023) explore the effects of constraining output in the text and feature domains. Taking VLP as the victim model, AdvDiffVLM (Guo et al., 2024) and AdvCLIP (Zhou et al., 2023a) employ Stable Diffusion (SD) (Rombach et al., 2022) and GAN (Goodfellow et al., 2020), respectively, to generate attack samples. IPI (Greshake et al., 2023) tries to disrupt model judgment by pasting text onto images in a way like watermarking (Jia et al., 2020). FigStep (Gong et al., 2023b), PPI (Chen et al., 2023b), and CAA (Shayegani et al., 2023) use typography to inject attack information into images, while MM-SB (Liu et al., 2023d) further integrates CIG images. Inspired by (Lan et al., 2019), TMM (Wang et al., 2024a) leverages Cross-Attention to identify regions where image and text features align, then strengthens transferability by replacing corresponding words or increasing perturbation weights.
Generalization. Unlike traditional attacks, which focus on Cross-Model, Cross-Image, and Cross-Environment attributes (see section 2.3), LVLM adversarial attacks introduce two new types of UAPs: Cross-prompt and Cross-corpus. Cross-prompt AEs induce incorrect responses under different text prompts, while cross-corpus AEs incorporate multiple references of malicious corpus. We collectively refer to cross-image/prompt/corpus samples as UAPs. Generalization in LVLMs can be improved through (Cross-Environment attributes needs further concerns):
@@ 7.4.2. Attack Purposes
As shown in Fig. 8 (a), we classify LVLM adversarial attacks into three types based on purposes:
Cognitive Bias. This attack manipulates LVLM to create cognitive distortions in its perception of text or images, leading to outputs that do not match the original input. For example, a cat may be misclassified as a dog, or an object in detection might be missed or mislocated. This type of attack can include tasks such as classification (Zhou et al., 2023a; Qraitem et al., 2024; Luo et al., 2024a; Cui et al., 2024; Greshake et al., 2023), detection (visual grounding) (Gao et al., 2024a; Wang et al., 2024a), image captioning (Chen et al., 2023c; Guo et al., 2024; Dong et al., 2023a; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023), VQA (Luo et al., 2024a; Cui et al., 2024; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023; Tu et al., 2023; Wang et al., 2024b), vision-language retrieval (Zhou et al., 2023a; Zheng et al., 2024; Wang et al., 2024a), and visual entailment (Wang et al., 2024a).
@@
Jailbreak. This attack uses carefully crafted inputs to trick the model into answering prohibited malicious questions correctly, bypassing safety alignment and avoiding refusal (OpenAI, 2024; AI, 2024; Yu et al., 2024). Jailbreak methods vary widely. To enhance the adversariality, we can add adversarial prefixes (Chen et al., 2023b) (e.g., designed through role-playing (Ma et al., 2024), context simulation (Wu et al., 2023), and RS/AA (Wu et al., 2023)) and suffixes (Wang et al., 2024e) (e.g., generated by GCG (Zou et al., 2023)) to the prompt, or align images with malicious text (Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024). Paired with textual pointers (e.g., using pronouns like “the objects” to refer to malicious content embedded in the image (Li et al., 2024c; Shayegani et al., 2023)), harmful information can be transferred from text to images through typography or CIG, forming an effective multimodal attack (Li et al., 2024c; Ma et al., 2024; Gong et al., 2023b; Shayegani et al., 2023; Liu et al., 2023d). RS/AA may also appears as aligned corpora in attacks (Ying et al., 2024b) and some tricks in LLM jailbreaks, like concealing malicious words with multilingual (Deng et al., 2023b) or obscure expressions (Xu et al., 2023b), remain untouched in LVLM.
Cognitive bias causes the model to err on “yes or no” questions, while prompt injection and jailbreak lead to mistakes on “right or wrong” questions. The boundary between jailbreak and prompt injection is not clear-cut; the main distinction lies in the attack objective, not the input type. Jailbreak aims to bypass safety mechanisms to output harmful content, while prompt injection seeks to hijack the model’s expected output for various attack goals. Jailbreak is essentially a subset of prompt injection, as bypassing safety mechanisms can also be a model hijacking goal.
@@
In fact, prompt injection is the most general concept of the three categories. If model hijacking results in the model generating prohibited malicious content, prompt injection manifests as a jailbreak attack. If the hijacking results in the model returning responses that differ from the original data, prompt injection displays cognitive bias.
Attack Applications. In addition to the three types of attack mentioned above, LVLM adversarial attacks have other motives and applications. For example, SRA (Wang et al., 2024b) focuses on disrupting the LVLM’s CoT process, thereby interfering with the reasoning process. Noting that LVLM deployment requires significant computational resources, VIA (Gao et al., 2024b) generates samples that force the victim model to output unnecessarily lengthy responses, maliciously consuming resources. Unlike other methods targeting a single model, TWW (Tan et al., 2024) and AgSmi (Gu et al., 2024) use collaboration or proxy networks formed by LVLMs to execute jailbreaks, discovering that malicious information can propagate between LVLM nodes with a high speed. AAMA (Wu et al., 2024) targets LVLM agents (a downstream application), interfering with the agent’s ability to assist users in classifieds, Reddit, and shopping scenarios.
@@ 7.4.3. Attack Techniques
In addition to the attacker’s knowledge and purposes, we have also summarized the common categories of techniques used in existing LVLM adversarial attacks. As shown in Fig. 8 (b), there are four commonly used techniques in LVLM adversarial attacks:
@@
These techniques are independent of each other; they can be used individually or in combination. When prompt manipulation is combined with other techniques, it forms a multimodal attack; otherwise, it remains a unimodal ones. In Fig. 8 (b), the multimodal attack (marked with an asterisk) represents the area where the Prompt Manipulation part (in blue) overlaps with other circles, while the unimodal attacks denotes the others.
@@
Unimodal Attacks. As shown in the non-overlapping green area in the lower left corner of Fig. 8 (b), this set of methods borrows from traditional adversarial attack approaches, where attack samples are constructed by applying perturbations to images through specific objective functions. By exploiting system prompts exposed due to vulnerabilities in GPT-4V, SASP (Wu et al., 2023) allows the victim LVLM to generate jailbreak prompts by itself, which is further enhanced by context simulation and RS/AA to increase success rates. PPI (Chen et al., 2023b), SGTA (Qraitem et al., 2024), and IPI (Greshake et al., 2023) use typography to inject misleading information into images, either jailbreaking the LVLM (Chen et al., 2023b) or disrupting its judgments (Qraitem et al., 2024; Greshake et al., 2023). InstructTA (Wang et al., 2023b) generates attack samples by constraining the distance between the target image and the perturbed image in the feature domain, where the target image is derived from CIG applied on the target text. Unimodal attacks may not fully exploit the attack surface of LVLMs, potentially limiting the desired impact.
@@
Multimodal Attacks. Multimodal attacks modify both the text and image inputs simultaneously, making fuller use of the LVLM’s attack surface. On the image side, they typically sample one or more from techniques such as Adversarial Perturbation, Typography, and CIG to construct effective adversarial images, while on the text side, the following three strategies are commonly used:
@@
Specifically, by combining manually designed text pointers and typographic images containing malicious information, FigStep (Gong et al., 2023b) and CAA (Shayegani et al., 2023) successfully carried out multimodal jailbreak attacks on models like GPT-4V and LLaVA. HADES (Li et al., 2024c), MM-SB (Liu et al., 2023d), and VRP (Ma et al., 2024) further incorporated CIG techniques to enhance the toxicity of attack samples. These methods first extract keywords or descriptions from harmful prompts through language models, then use Typography and CIG to generate different harmful images. After concatenating the images, they combine them with the modified prompt (Gong et al., 2023b; Shayegani et al., 2023; Li et al., 2024c; Liu et al., 2023d) or simulate a role-play scenario (Ma et al., 2024) to launch the attack. AnyDoor (Lu et al., 2024) binds cross-image perturbations with text triggers to create backdoor attacks. UMK (Wang et al., 2024e) and BAP (Ying et al., 2024b) first generate cross-corpus images, then use GCG (Zou et al., 2023) to generate adversarial suffixes or let the victim LVLM generate effective text prompts by itself. Aware of the importance of modality-consistent features for transferability, TMM (Wang et al., 2024a) enhances transferability by replacing words and increasing image perturbations in the regions that text and image aligned.

### v2

@@ 7.3.3. Metrics
Here, represents the set of attack samples, and is a binary function used to determine whether an attack sample is effective. Since the output of an LVLM is natural text, the criteria for the function can vary across different attack scenarios. For example, in image captioning tasks, the determination of effectiveness is often made by checking whether the output text diverges from the original meanings (untargeted) (Dong et al., 2023a) or close to the target semantics (targeted) (Wang et al., 2023b). In prompt injection and jailbreak attacks, success is usually determined by scrutinizing whether the output contains malicious contents (Lu et al., 2024; Fu et al., 2023) or violates usage policies (Chen et al., 2024b; Ying et al., 2024b). As summarized in Fig. 9, the existing implementations of the function in research can be categorized into four types:
[…]
@@
String Matching. This method checks whether the LVLM output contains certain keywords/phrases in the focus list or exactly matches a predefined target text (Lu et al., 2024; Fu et al., 2023). It is the most convenient strategy but can lead to false positives or negatives in some cases (e.g., when certain keywords are not in the focus list or the list is incomplete). In addition, the method is also known as Contain and ExactMatch (Liu et al., 2024a).
@@ 7.4.1. Attacker Knowledge
As shown in Fig. 10 (a), in LVLM adversarial attacks, we follow traditional criteria and categorize knowledge access into white-box, gray-box, and black-box. Unlike previous categories, we place query-based methods that require interaction with the victim model under the gray-box category, as these methods gather some degree of information in queries (e.g., estimating gradients in the text feature domain by LVLM’s responses (Zhao et al., 2024) or to generate/update attack prompts (Qraitem et al., 2024; Yin et al., 2023; Ma et al., 2024)), making them not entirely ignorant of the victims. Black-box methods refer to transfer-based attacks, which do not seek any information from the victims. These methods rely solely on prior knowledge, such as publicly available information and educated guesses. section 4 discusses the detailed classification criteria.
@@
Specifically, QT-Atk (Zhao et al., 2024) draws on the cascading approach of traditional transfer and query-based attacks (Cheng et al., 2019; Dong et al., 2021). It first applies transfer-based attacks to compute preliminary image perturbations and then refines the results (with PGD (Madry, 2017)) by estimating gradients using finite differences based on the model’s responses. All of (Qraitem et al., 2024; Yin et al., 2023; Ma et al., 2024) adopt the self-generate approach: the difference is that SGTA (Qraitem et al., 2024) requires identifying the most confusing class and description for a given image, while SASP (Yin et al., 2023) directly updates the attack prompt and VRP (Ma et al., 2024) incorporates role-playing dialogue. InstructTA (Wang et al., 2023b), on the other hand, assumes access to the victim’s visual encoder and applies a feature attack on it.
@@
Specifically, HADES (Li et al., 2024b) employs a three-stage strategy, progressively adding harmful information through typography, CIG, and adversarial perturbations to the attack samples. Focusing on disrupting LVLM’s CoT (Wei et al., 2022) process, SRA (Wang et al., 2024d) attacks both the reasoning and answer generation components. BMA (Bailey et al., 2023), JBA (Qi et al., 2024), UMK (Wang et al., 2024e), and BAP (Ying et al., 2024a) aggregate perturbations across datasets to create cross-prompt (Bailey et al., 2023)/corpus (Qi et al., 2024; Wang et al., 2024e; Ying et al., 2024a) universal adversarial images. AnyDoor (Lu et al., 2024) takes this further by binding universal adversarial images to specific text triggers (e.g., “SUDO”) to execute backdoor attacks. Unlike previously aggregated universal samples, CroPA (Luo et al., 2024a) achieves cross-prompt effects by continuously perturbing input text during sample “training”, which achieves aggregation over perturbed input text.
@@
Constructing attacks based on traditional black-box methods in different ways, including directly applying (Chen et al., 2023d; Dong et al., 2023a), with certain adjustments (Niu et al., 2024; Wu et al., 2024; Tu et al., 2023; Li et al., 2025), and by generative models (Guo et al., 2024; Zhou et al., 2023a).
Leveraging techniques such as typography and CIG, combined with specially designed text, to create attacks specific to LVLMs (Greshake et al., 2023; Gong et al., 2023b; Chen et al., 2023c; Shayegani et al., 2023; Liu et al., 2023d). (This type of methods is similar to that of white-box attacks, but researchers report the cross-model transferability on it.)
@@
Specifically, CWA (Chen et al., 2023d) and SSA-CWA (Dong et al., 2023a) directly apply traditional transfer attack methods to test LVLMs, while M-Attack (Li et al., 2025), ImgJP (Niu et al., 2024), AAMA (Wu et al., 2024), and OOD-VQA (Tu et al., 2023) explore the effects of constraining output in the text and feature domains. Taking VLP as the victim model, AdvDiffVLM (Guo et al., 2024) and AdvCLIP (Zhou et al., 2023a) employ Stable Diffusion (SD) (Rombach et al., 2022) and GAN (Goodfellow et al., 2020), respectively, to generate attack samples. IPI (Greshake et al., 2023) tries to disrupt model judgment by pasting text onto images in a way like watermarking (Jia et al., 2020). FigStep (Gong et al., 2023b), PPI (Chen et al., 2023c), and CAA (Shayegani et al., 2023) use typography to inject attack information into images, while MM-SB (Liu et al., 2023d) further integrates CIG images. Inspired by (Lan et al., 2019), TMM (Wang et al., 2024a) leverages Cross-Attention to identify regions where image and text features align, then strengthens transferability by replacing corresponding words or increasing perturbation weights.
Generalization. Unlike traditional attacks, which focus on Cross-Model, Cross-Image, and Cross-Environment attributes (see section 2.3), LVLM adversarial attacks introduce two new types of UAPs: Cross-prompt and Cross-corpus. Cross-prompt AEs induce incorrect responses under different text prompts, while cross-corpus AEs incorporate multiple references of malicious corpus. We collectively refer to cross-image/prompt/corpus samples as UAPs. Generalization in LVLMs can be improved through (Cross-Environment attributes need further concerns):
@@ 7.4.2. Attack Purposes
Cognitive Bias. This attack induce cognitive distortions in LVLMs, generating outputs that misinterpret text or images. For example, mistaking a cat for a dog, or mislocating the objects. It impacts tasks like classification (Zhou et al., 2023a; Qraitem et al., 2024; Luo et al., 2024a; Cui et al., 2024; Greshake et al., 2023), detection (visual grounding) (Gao et al., 2024a; Wang et al., 2024a), image captioning (Chen et al., 2023d; Guo et al., 2024; Dong et al., 2023a; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023), VQA (Luo et al., 2024a; Cui et al., 2024; Zhao et al., 2024; Wang et al., 2023b; Schlarmann and Hein, 2023; Tu et al., 2023; Wang et al., 2024d), vision-language retrieval (Zhou et al., 2023a; Zheng et al., 2024; Wang et al., 2024a), and visual entailment (Wang et al., 2024a).
@@
Jailbreak. This attack uses carefully crafted inputs to trick the model into answering prohibited malicious questions correctly, bypassing safety alignment and avoiding refusal (OpenAI, 2024; AI, 2024; Yu et al., 2024). Jailbreak methods vary widely. To enhance the adversariality, we can add adversarial prefixes (Chen et al., 2023c) (e.g., designed through role-playing (Ma et al., 2024), context simulation (Yin et al., 2023), and RS/AA (Yin et al., 2023)) and suffixes (Wang et al., 2024e) (e.g., generated by GCG (Zou et al., 2023)) to the prompt, or align images with malicious text (Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024). Paired with textual pointers (e.g., using pronouns like “the objects” to refer to malicious content embedded in the image (Li et al., 2024b; Shayegani et al., 2023)), harmful information can be transferred from text to images through typography or CIG, forming an effective multimodal attack (Li et al., 2024b; Ma et al., 2024; Gong et al., 2023b; Shayegani et al., 2023; Liu et al., 2023d). RS/AA may also appears as aligned corpora in attacks (Ying et al., 2024a) and some tricks in LLM jailbreaks, like concealing malicious words with multilingual (Deng et al., 2023b) or obscure expressions (Xu et al., 2023b), remain untouched in LVLM.
Cognitive bias leads to errors in “yes or no”, while prompt injection and jailbreak lead to mistakes on “right or wrong”. The boundary between these two is not clear-cut; the main distinction lies in the attack objective, not the input type. Jailbreak aims to bypass safeguards to produce harmful content, while prompt injection seeks to hijack models’ output. Jailbreak is essentially a subset of prompt injection, as bypassing safeguards can be a form of model hijacking.
@@
In fact, prompt injection is the broadest concept. If model hijacking leads to malicious content generation, it acts as a jailbreak. If responses deviate from the original data, it introduces cognitive bias.
Attack Applications. LVLM adversarial attacks have diverse motives and applications beyond the three mentioned above. SRA (Wang et al., 2024d) disrupts the LVLM’s CoT process, affecting reasoning. VIA (Gao et al., 2024b) generates AEs that force excessive responses, draining computational resources. TWW (Tan et al., 2024) and AgSmi (Gu et al., 2024) exploit collaborative LVLM networks, demonstrating rapid propagation of malicious information across nodes during jailbreak attempts. AAMA (Wu et al., 2024) targets LVLM agents (a downstream application), impairing their assistance in classifieds, Reddit, and shopping scenarios.
@@ 7.4.3. Attack Techniques
Illustration of Attack Purposes and Techniques.
Besides the attacker’s knowledge and intent, attack techniques are equally critical. Fig. 10 (b) illustrates a technology-based taxonomy, while Fig. 11 (b) provides a simplified schematic:
@@
These techniques can be used alone or combined. When prompt manipulation is paired with others, it forms a multimodal attack; otherwise, it remains unimodal. In Fig. 10 (b), multimodal attacks are asterisked marked.
@@
The non-overlapping green region in Fig. 10 (b) illustrates methods drawing on traditional adversarial attacks, where images are perturbed via specific objective functions. SASP (Yin et al., 2023) exploits GPT-4V vulnerabilities to make LVLMs generate jailbreak prompts by itself, enhanced by context simulation and RS/AA for higher success rates. PPI (Chen et al., 2023c), SGTA (Qraitem et al., 2024), and IPI (Greshake et al., 2023) use typography to inject misleading information into images, either jailbreaking LVLMs (Chen et al., 2023c) or disrupting their judgments (Qraitem et al., 2024; Greshake et al., 2023). InstructTA (Wang et al., 2023b) generate AEs by constraining feature-domain distances between target and perturbed images, with target images derived via CIG based on target texts. Since unimodal attacks may not fully exploit the attack surface of LVLMs, their impact could be limited.
@@
Multimodal Attacks. Multimodal attacks manipulate both text and image inputs, exploiting fuller attack surface of LVLMs. They generate adversarial images using one or more techniques like Adversarial Perturbation, Typography, and CIG, while employing three common strategies to modify text:
@@
Specifically, by combining manually designed text pointers and typographic images containing malicious information, FigStep (Gong et al., 2023b) and CAA (Shayegani et al., 2023) successfully carried out multimodal jailbreak attacks on models like GPT-4V and LLaVA. HADES (Li et al., 2024b), MM-SB (Liu et al., 2023d), and VRP (Ma et al., 2024) further incorporated CIG techniques to enhance the toxicity of attack samples. These methods first extract keywords or descriptions from harmful prompts through language models, then use Typography and CIG to generate different harmful images. After concatenating the images, they combine them with the modified prompt (Gong et al., 2023b; Shayegani et al., 2023; Li et al., 2024b; Liu et al., 2023d) or simulate a role-play scenario (Ma et al., 2024) to launch the attack. AnyDoor (Lu et al., 2024) binds cross-image perturbations with text triggers to create backdoor attacks. UMK (Wang et al., 2024e) and BAP (Ying et al., 2024a) first generate cross-corpus images, then use GCG (Zou et al., 2023) to generate adversarial suffixes or let the victim LVLM generate effective text prompts by itself. Aware of the importance of modality-consistent features for transferability, TMM (Wang et al., 2024a) enhances transferability by replacing words and increasing image perturbations in the regions that text and image aligned.